OpenStudy (anonymous):

how do you factor 15f^2-31fg+10g^2

OpenStudy (anonymous):

Alright this is how you go about doing this.. you have letters and numbers here to worry about.. since you have f's and g's, you have a lot of factors here... You know you need to FOIL 2 equations to come up with this answer... Thus you will have (?f +/- ?g) (?f +/- ?g), now you need to figure out your symbols... you know you have to multiply 2 numbers to get a + value but add em up to get a - value. Thus you can assume both numbers have to be negative... thus you will have (? f - ? g) (? f - ? g) Now you just need your numbers... I could just give you the answers, but try it out for yourself first.. if you need additional help just let me know.

OpenStudy (anonymous):

i dont get it . please help

OpenStudy (anonymous):

Alright well you need factors that multiply to 15 for the f's so what 2 numbers can multiply to get 15?

OpenStudy (anonymous):

3 and 5

OpenStudy (anonymous):

alright so put those numbers in.. so you will have (3f - ? g) (5f - ? g). You know how to FOIL correct?

OpenStudy (anonymous):

okay then what

OpenStudy (anonymous):

you need to get 2 numbers that multiply to 10, but also add up to 31.. keep in mind the the 3 and 5 you just picked will play a factor too... but until then.. what 2 numbers multiply to 10?

OpenStudy (anonymous):

2 and 5

OpenStudy (anonymous):

ok so now you need to put 2 and 5 into the equation.. but which one goes first? A) (3f - 5g) (5f - 2g) B) (3f - 2g) (5f - 5g)

OpenStudy (anonymous):

b right ?

OpenStudy (anonymous):

you're just guessing.. Foil out B and tell me what you get.

OpenStudy (anonymous):

i dont know what that means , i just need to figure out the answer and how to get it

OpenStudy (anonymous):

FOIL is a way to solve 2 equations.. it means.. First, Outside, Inside, Last Let me give you an example first... Example: (x +2 ) (x + 3) First (F): x(x) = x^2 Outside (O): x(3) = +3x Inside (I): x(2) = +2x Last (L): (3)2 = +6 Thus you will get x^2 +3x + 2x +6... and you would combine them to get x^2+5x+6 Now try it with the equation above.....

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

Does it help or giving up?

OpenStudy (anonymous):

giving up

OpenStudy (anonymous):

Don't give up or you will continue to struggle....

OpenStudy (anonymous):

I'd give you the answer but then you wouldn't be learning anything. Just try to FOIL it and if you don't get it from there I can help you